OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = –6 3x – 2y = 11 What is the solution to the system?

OpenStudy (anonymous):

Answer possibilities are: (2, –1) (1, –4) (3, –2) (–1, –7) Last time I tried to solve it, I got 2, -1, and it was marked incorrect. But that's always the answer I get. :/ So.. I must be doing something wrong. x.x Help, please? :o

jimthompson5910 (jim_thompson5910):

Use a graphing calculator to get what you see attached You'll see that the two lines intersect at the point (1,-4) So the answer is choice B

jimthompson5910 (jim_thompson5910):

Oh I should point out that the green line is 2x + 2y = –6 and the blue line is 3x – 2y = 11

OpenStudy (anonymous):

Ohh. I know it said to solve by graphing, but I tried to solve by substitution instead. :/ But thank you. :D

jimthompson5910 (jim_thompson5910):

If you want to use substitution, then... Solve the first equation for y 2x + 2y = –6 2y = –6 - 2x y = -6/2 - 2x/2 y = -3 - x -------------------------- Then plug this into the second equation 3x – 2y = 11 3x – 2(-3 - x) = 11 3x + 6 + 2x = 11 5x + 6 = 11 5x = 11 - 6 5x = 5 x = 5/5 x = 1 <---- First part of the answer --------------------------------- Now use this to find y y = -3 - x y = -3 - 1 y = -4 <---- Second part of the answer So the solutions are x = 1 and y = -4 giving us the ordered pair (1, -4)

OpenStudy (anonymous):

OH I meant to say elimination o.o

OpenStudy (anonymous):

Either way, I probably did it wrong. xD But thank you very much. n.n

jimthompson5910 (jim_thompson5910):

Add the two equations 2x + 2y = –6 3x – 2y = 11 ------------ 5x + 0y = 5 which gives us 5x = 5 Solve for x to get x = 1 Then use this to find y 3x – 2y = 11 3(1) – 2y = 11 3 - 2y = 11 -2y = 11 - 3 -2y = 8 y = 8/(-2) y = -4 So we get the same answer

OpenStudy (anonymous):

Oh, I see what I did wrong now. xD

jimthompson5910 (jim_thompson5910):

that's great :)

OpenStudy (anonymous):

can you break it down ?